### v2

# thanks: Corresponding author

Abstract
We study the exotic $\Delta^{0}\Delta^{0}$ dibaryon state with quantum numbers $J^{P}=2^{+}$ within the framework of QCD sum rules. A tensor interpolating current is constructed to determine the mass and residue of this state, including explicit calculations of contributions from quark, gluon, and mixed vacuum condensates up to dimension nine. The numerical analysis yields a mass of $m_{\Delta^{0}\Delta^{0}}=2426^{+101}_{-108}~\mathrm{MeV}$ and a residue of $f_{\Delta^{0}\Delta^{0}}=\left(2.30^{+0.49}_{-0.45}\right)\times 10^{-4}~\mathrm{GeV}^{8}$. The predicted mass lies about $38~\mathrm{MeV}$ below the $2m_{\Delta^{0}}$ threshold, which indicates the possible existence of a bound configuration. Consequently, the tensor $\Delta^{0}\Delta^{0}$ dibaryon may serve as a promising candidate for future experimental searches.
I Introduction
The investigation of multiquark states within quantum chromodynamics (QCD) began shortly after the theory was established in the 1970s. A pioneering bag-model calculation from that period ignited interest in these exotic systems and motivated both experimental searches and alternative theoretical approaches. While conventional hadrons, such as mesons (quark–antiquark pairs) and baryons (three-quark states), were well understood within the quark model, the possibility of exotic hadrons, such as tetraquarks (four quarks) and pentaquarks (five quarks), attracted considerable theoretical attention Gell-Mann:1964ewy . Early theoretical models predicted the existence of multiquark states, but experimental confirmation remained elusive for decades.
Significant progress occurred in the 2000s due to advancements in experimental techniques and particle accelerators. In 2003, the Belle Collaboration reported the discovery of the $X(3872)$, widely regarded as the first strong tetraquark candidate. This was followed by further discoveries, including the observation of pentaquark states by the LHCb experiment at CERN, where so-called “exotic” hidden-charm XYZXY\!Z tetraquark and $P_{c}$ pentaquark states were reported. These milestones marked a turning point in multiquark research, opening new avenues for exploring QCD and deepening our understanding of hadron structure and the strong interactions that govern their behavior Belle:2003nnu ; Swanson:2006st ; LHCb:2015yax ; Chen:2016qju ; Lebed:2016hpi ; Olsen:2017bmm .
Dibaryons, as another important class of multiquark hadrons, pose some of the most complex theoretical and experimental challenges in strong interaction physics. The deuteron, discovered in 1932 by Urey, Brickwedde, and Murphy Urey:1932gik , was the first identified dibaryon. Another intriguing example is the H-dibaryon predicted by Jaffe Jaffe:1976yi , which continues to attract both experimental and theoretical interest. More recently, the $d^{*}(2380)$ resonance was observed in the total cross section at $\sqrt{s}=2.37~\mathrm{GeV}$ with $\Gamma\approx 70~\mathrm{MeV}$ and quantum numbers $I(J^{P})=0(3^{+})$, in double-pionic fusion reactions reported by the WASA-at-COSY Collaboration WASA-at-COSY:2011bjg ; WASA-at-COSY:2012seb ; Gal:2013dca ; WASA-at-COSY:2014dmv ; WASA-at-COSY:2014qkg . Following its observation, numerous studies have examined the $d^{*}(2380)$, including investigations of its total width Dong:2015cxa , as well as analyses of its mass, binding energy, root-mean-square radius, and electromagnetic form factors in comparison with experimental data Dong:2018ryf ; Dong:2019zvg . The $d^{*}(2380)$ has also been studied using QCD sum rules Chen:2014vha , where three $\Delta$$\Delta$-like interpolating currents yielded a predicted mass $M_{d^{*}}=2.4\pm 0.2~\mathrm{GeV}$, consistent with the WASA measurement. Further research has investigated its structure and decay properties using a chiral constituent quark model Huang:2019lzt , evaluated the properties of its decuplet states Bashkanov:2020lxv , and explored its possible configurations as a hexaquark state Kim:2020rwn .
Recent investigations have examined the influence of the $d^{*}(2380)$ hexaquark on the equation of state (EoS) of dense neutron star matter, exploring the implications of its existence for neutron star properties. These studies have shown that including the $d^{*}$ state notably softens the EoS at high densities, thereby reducing the predicted maximum mass of neutron stars Celi:2023gtj ; Celi:2025wnc . Motivated by the possibility of a stable hexaquark $(uuddss)$ as a dark matter candidate, Ref. Shahrbaf:2022upc investigated its compatibility with neutron star observations, focusing on constraints from maximum mass and tidal deformability.
In parallel, interest in dibaryons has been fueled by proposals to explore them as potential dark matter candidates Farrar:2003qy ; Farrar:2017eqq ; Farrar:2018hac . Studies on the mass and stability of the H-dibaryon, inspired by Jaffe’s original prediction Jaffe:1976yi , have been carried out using various particle physics models. The corrected MIT bag model yields $m_{H}=2240~\mathrm{MeV}$, placing it just above the $2m_{\Lambda}$ threshold Halprin:1982pb , while a chiral model predicts a substantially lower value of $1130~\mathrm{MeV}$ Yost:1985mj .
[…]
The interaction between the nucleon ($N$) and the Omega baryon ($\Omega$) was investigated in Ref. HALQCD:2014okw , where a bound state was identified with a binding energy of $18.9(5.0)^{+12.1}_{-1.8}~\mathrm{MeV}$. Lattice QCD studies of the $\Omega\Omega$ dibaryon in the $^{1}S_{0}$ channel reported a binding energy of $B_{\Omega\Omega}=1.6(6)^{+0.7}_{-0.6}~\mathrm{MeV}$, indicating an overall attraction near the unitary regime Gongyo:2017fjb . Following this, analyses of two-particle momentum correlations for $\Omega\Omega$ and $p\Omega$ pairs in relativistic heavy-ion collisions predicted shallow bound states characterized by comparatively large positive scattering lengths in the $J=0$ $\Omega\Omega$ and $J=2$ $p\Omega$ channels Morita:2019rph .
Using QCD sum rules, the exotic $\Omega\Omega$ dibaryons with $J^{P}=0^{+}$ and $2^{+}$ have been studied in a molecular picture. The predicted scalar state has a mass of $(3.33\pm 0.22)~\mathrm{GeV}$, suggesting a loosely bound system, while the tensor state at $(3.24\pm 0.23)~\mathrm{GeV}$ may be more deeply bound Chen:2019vdh . The production of $p\Omega$ ($^{5}S_{2}$) and $\Omega\Omega$ ($^{1}S_{0}$) in central Pb+Pb collisions at $\sqrt{s_{NN}}=2.76~\mathrm{TeV}$ was studied in Ref. Pu:2024kfh , considering two possible structural scenarios: molecular states and compact six-quark states.
In Ref. Junnarkar:2024kwd , the ground state spectrum of two-flavor heavy quark baryons within the spin-singlet channel was investigated. Moreover, the relationship between the binding energy and the constituent quark masses was explored by varying quark masses across a range from strange to bottom quark masses. The S-wave $\Lambda\Lambda$ and $N\Xi$ systems were examined using lattice QCD HALQCD:2019wsz , with results indicating that the $N\Xi$ interaction lies close to unitarity, whereas the $\Lambda\Lambda$ interaction at low energies does not support a bound state.
[…]
In the heavy-flavor sector, studies of the mass spectra of molecular-type hexaquark states have indicated the possible existence of molecular configurations in the $^{*}, DDD¯D$$\bar{D}^{*}$, and $\Sigma_{c}\bar{D}^{(*)}$ systems, suggesting the emergence of numerous deuteron-like hexaquarks Wang:2024riu . In Ref. Geng:2024dpk , the study of doubly charmed $H$-like dibaryon $\Lambda_{c}\Lambda_{c}$ scattering from lattice QCD shows that the $\Lambda_{c}\Lambda_{c}$ system behaves as a typical scattering state with no bound state formation at the given quark mass. Furthermore, the investigation of doubly-charmed hexaquark configurations was conducted within the diquark framework, employing both the constituent quark model and the quark-interchange approach. The analysis, based on key properties such as mass spectra, suggests the nonexistence of stable bound states. Nonetheless, the evaluated root-mean-square radii indicate that these hexaquark systems exhibit a notably compact configuration An:2025rjv . The possibility of the existence of dibaryon states composed of heavy baryons, such as $\Xi^{*}_{c}\Xi^{*}_{c}$ and $\Xi_{c}\Xi_{c}$, has been investigated Lee:2011rka ; Shiri:2023fjh ; Shah:2024thr . In Ref. Shah:2024thr , calculations based on a variational method yield masses of $5.291~\mathrm{GeV}$ for the $J^{P}=0^{+}$ channel with total isospin $I=1$, and $5.281~\mathrm{GeV}$ for the $J^{P}=1^{+}$ channel with $I=0$. Fully heavy dibaryons represent a particularly intriguing and debated topic. In Ref. Richard:2020zxb , the authors explored the possibility of bbbcccbbbccc dibaryons and reported no bound states for either the $\Omega_{ccc}\Omega_{ccc}$ or $\Omega_{bbb}\Omega_{bbb}$ systems. Conversely, Ref. Lyu:2022tsd indicated that the $\Omega_{ccc}\Omega_{ccc}$ system is weakly bound with a binding energy of $5.68(0.77)~\mathrm{MeV}$, while Ref. Mathur:2022ovu predicted a deeply bound $\Omega_{bbb}\Omega_{bbb}$ state, with a binding energy of $81^{+14}_{-16}~\mathrm{MeV}$. More recently, Ref. Martin-Higueras:2024qaw conducted a renewed analysis of the potential existence of $\Omega_{ccc}\Omega_{ccc}$ and $\Omega_{bbb}\Omega_{bbb}$ dibaryons using a constituent quark model. Their findings suggest that six charm or six bottom quarks can form bound states with quantum numbers $J^{P}=0^{+}$, yielding binding energies of $0.71~\mathrm{MeV}$ for the charm case and $1.98~\mathrm{MeV}$ for the bottom case.
As outlined in the aforementioned studies, several theoretical frameworks have been utilized to investigate six-quark systems and dibaryons. These include QCD sum rules, lattice QCD, chiral perturbation theory, quark models, and effective field theories. Among these, the QCD sum rule approach has proven to be a powerful nonperturbative method, extensively applied to various hadronic phenomena, including hadron spectroscopy and decay processes. Within the QCD sum rule framework, suitable interpolating currents corresponding to the hadron of interest are constructed. These currents are employed to formulate two-point or three-point correlation functions, which are used to extract the hadron’s mass and decay properties, respectively. By matching the operator product expansion (OPE) of the correlation function to its hadronic spectral representation, one derives sum rules that enable the determination of physical observables.
One of the key advantages of the QCD sum rule method is that it is firmly grounded in the fundamental QCD Lagrangian. The formalism is fully relativistic and free from arbitrary or purely phenomenological parameters, thereby ensuring a high degree of theoretical consistency and predictive reliability.
In the light quark sector, the mass and coupling constant of a scalar six-quark state composed of uuddssuuddss quarks have been studied using the QCD sum rule approach to explore its viability as a dark matter candidate Azizi:2019xla . The possibility of a $\Delta^{0}\Delta^{0}$ dibaryon has also been investigated through various methods, such as the quark delocalization color screening model and the chiral quark model Huang:2013nba . More recently, the scalar $\Delta^{0}\Delta^{0}$ state was studied using QCD sum rules Mutuk:2022zgn , where the mass was estimated as $m_{D}=2326^{+114}_{-126}~\mathrm{MeV}$, and the decay constant was found to be approximately $f_{D}=\left(2.94^{+0.30}_{-0.34}\right)\times 10^{-4}~\mathrm{GeV}^{7}$. In the present work, we extend this line of investigation by calculating the mass and decay constant (residue) of a hypothetical $\Delta^{0}\Delta^{0}$ dibaryon in the tensor state with quantum numbers $J^{P}=2^{+}$ using the QCD sum rule method.
II METHODOLOGY
To investigate dibaryon systems within the framework of QCD sum rules, the $\Delta^{0}\Delta^{0}$ interpolating currents are constructed using the local Ioffe current representation of the $\Delta^{0}$ baryon:
$\eta_{\mu}=\frac{1}{\sqrt{3}}\,\varepsilon^{abc}\left[2\left(d_{a}^{T}C\gamma_{\mu}u_{b}\right)d_{c}+\left(d_{a}^{T}C\gamma_{\mu}d_{b}\right)u_{c}\right],$ (1)
where $C=i\gamma_{2}\gamma_{0}$ is the charge conjugation matrix, $a$, $b$, $c$ denote color indices, $\gamma_{\mu}$ are Dirac gamma matrices, and $T$ indicates the transpose operator.
Using this interpolating current, the $\Delta^{0}\Delta^{0}$ dibaryon with quantum numbers $J^{P}=2^{+}$ is constructed as
$J_{\mu\nu}^{\Delta^{0}\Delta^{0}}=\frac{1}{3}\,\epsilon^{abc}\epsilon^{def}\left[2\left(d^{a\,T}C\gamma_{\mu}u^{b}\right)d^{c}+\left(d^{a\,T}C\gamma_{\mu}d^{b}\right)u^{c}\right]\cdot C\gamma_{5}\cdot\left[2d^{d\,T}\left(d^{e}C\gamma_{\nu}u^{f}\right)+u^{d\,T}\left(d^{e}C\gamma_{\nu}d^{f}\right)\right].$ (2)
To extract physical properties such as the mass and residue, we employ the two-point QCD sum rule method by analyzing the correlation function
$\Pi_{\mu\nu\,\rho\sigma}(p^{2})=i\int d^{4}x\,e^{ip\cdot x}\langle 0|T\{J^{\Delta^{0}\Delta^{0}}_{\mu\nu}(x)J^{\Delta^{0}\Delta^{0}\dagger}_{\rho\sigma}(0)\}|0\rangle,$ (3)
where $T$ is the time-order operator and $J_{\mu\nu(\rho\sigma)}^{\Delta^{0}\Delta^{0}}$ is the interpolating current for the tensor dibaryon state. In the framework of QCD sum rules, the correlation function is evaluated from two complementary perspectives. On the physical (phenomenological) side, the correlation function is modeled in terms of hadronic degrees of freedom in the time-like region, allowing for the extraction of physical parameters such as the mass and residue in different states. On the QCD (theoretical) side, the same correlation function is computed in the space-like region using quark and gluon fields. This representation includes contributions from the QCD coupling constant, quark masses, and gluon condensates of various dimensions, capturing the non-perturbative aspects of the QCD vacuum. By equating the two sides through dispersion relations and invoking the quark-hadron duality assumption, one can extract the hadronic parameters in terms of QCD quantities. To enhance the reliability of the sum rule and suppress contributions from higher resonances and continuum states, Borel transformation and continuum subtraction procedures are employed. We commence by evaluating the hadronic (physical) representation of the correlation function through the insertion of appropriate complete sets of states, possessing the same quantum numbers as the interpolating current, into the corresponding positions. A complete set of intermediate hadronic states is inserted into the correlation function presented in Eq. (3). After performing the integration over $x$, the correlation function can be written as:
$\Pi^{\text{Phys}}_{\mu\nu\rho\sigma}(p)=\frac{\langle 0|j_{\mu\nu}(0)|D(p)\rangle\langle D(p)|J_{\rho\sigma}(0)|0\rangle}{m_{D}^{2}-p^{2}}+\cdots,$ (4)
where $\cdots$ represents the higher states and continuum contributions. The matrix elements responsible for creating hadronic states from the vacuum can be parameterized in terms of the decay constant and the corresponding polarization tensor as follows:
$\langle 0|J^{\Delta^{0}\Delta^{0}}_{\mu\nu}|D(p)\rangle={f_{\Delta^{0}\Delta^{0}}}\varepsilon(p)^{\lambda}_{\mu\nu}+\cdots.$ (5)
Here, $f_{\Delta^{0}\Delta^{0}}$ denotes the coupling constant, and $\varepsilon_{\mu\nu}$ is the polarization tensor associated with the spin-2 state. The polarization sum is given by
$\sum_{\lambda}\varepsilon^{(\lambda)}_{\mu\nu}\varepsilon^{*(\lambda)}_{\rho\sigma}=\frac{1}{2}\left(T_{\mu\rho}T_{\nu\sigma}+T_{\mu\rho}T_{\nu\rho}\right)-\frac{1}{3}T_{\mu\nu}T_{\rho\sigma},$ (6)
where
$T_{\mu\nu}=-g_{\mu\nu}+\frac{p_{\mu}p_{\nu}}{p^{2}}.$ (7)
Combining the above results and summing over polarization states via the corresponding completeness relation yields the physical representation of the correlation function in the following final form
$\Pi^{\text{Phys}}_{\mu\nu\rho\sigma}(p)=\frac{f_{\Delta^{0}\Delta^{0}}^{2}}{m_{D}^{2}-p^{2}}\left[\frac{1}{2}\left(g_{\mu\rho}g_{\nu\sigma}+g_{\mu\sigma}g_{\nu\rho}\right)\right]+\text{other structures}+\cdots,$ (8)
where the ellipses represent contributions from additional Lorentz structures, higher excited, and continuum states. The term proportional to $(g_{\mu\rho}g_{\nu\sigma}+g_{\mu\sigma}g_{\nu\rho})$ corresponds solely to the contribution from the spin-2 particle, while the remaining terms arise from the presence of spin-0 and spin-1 states as well.
In the next step of our analysis, we evaluate the correlator $\Pi_{\mu\nu\rho\sigma}(p)$ using OPE. For this purpose, we substitute the explicit form of the current $J_{\mu\nu}(x)$ into Eq. (3) and perform the necessary contractions of the quark fields via Wick’s theorem to derive $\Pi_{\mu\nu\rho\sigma}^{\mathrm{QCD}}(p)$. This leads to
$\Pi^{\text{QCD}}_{\mu\nu\rho\sigma}(p)=\frac{16}{9}\int d^{4}x\,e^{ip\cdot x}\times\Big($ (9)
$\delta_{af}\delta_{be}\delta_{cd}\,\delta_{a^{\prime}f^{\prime}}\delta_{b^{\prime}e^{\prime}}\delta_{c^{\prime}d^{\prime}}\,\operatorname{Tr}\big(S^{d}_{ad^{\prime}}(x,0)\gamma^{\rho}\tilde{S}^{u}_{be^{\prime}}(x,0)\gamma^{\mu}\big)\operatorname{Tr}\big(S^{d}_{cc^{\prime}}(x,0)\gamma^{5}\tilde{S}^{d}_{ff^{\prime}}(x,0)\gamma^{5}\big)$
$\qquad\times\operatorname{Tr}\big(S^{d}_{da^{\prime}}(x,0)\gamma^{\sigma}\tilde{S}^{u}_{eb^{\prime}}(x,0)\gamma^{\nu}\big)$
- $\delta_{ae}\delta_{bf}\delta_{cd}\,\delta_{a^{\prime}f^{\prime}}\delta_{b^{\prime}e^{\prime}}\delta_{c^{\prime}d^{\prime}}\,\operatorname{Tr}\big(S^{d}_{ad^{\prime}}(x,0)\gamma^{\rho}\tilde{S}^{u}_{be^{\prime}}(x,0)\gamma^{\mu}\big)\operatorname{Tr}\big(S^{d}_{cc^{\prime}}(x,0)\gamma^{5}\tilde{S}^{d}_{ff^{\prime}}(x,0)\gamma^{5}\big)$
$\qquad\times\operatorname{Tr}\big(S^{d}_{da^{\prime}}(x,0)\gamma^{\sigma}\tilde{S}^{u}_{eb^{\prime}}(x,0)\gamma^{\nu}\big)$
- $\delta_{af}\delta_{bd}\delta_{ce}\,\delta_{a^{\prime}d^{\prime}}\delta_{b^{\prime}f^{\prime}}\delta_{c^{\prime}e^{\prime}}\,\operatorname{Tr}\big(S^{d}_{ad^{\prime}}(x,0)\gamma^{\rho}\tilde{S}^{u}_{be^{\prime}}(x,0)\gamma^{\mu}\big)\operatorname{Tr}\big(S^{d}_{cc^{\prime}}(x,0)\gamma^{5}\tilde{S}^{d}_{ff^{\prime}}(x,0)\gamma^{5}\big)$
$\qquad\times\operatorname{Tr}\big(S^{d}_{da^{\prime}}(x,0)\gamma^{\sigma}\tilde{S}^{u}_{eb^{\prime}}(x,0)\gamma^{\nu}\big)$
$+$ $\delta_{ad}\delta_{bf}\delta_{ce}\,\delta_{a^{\prime}e^{\prime}}\delta_{b^{\prime}d^{\prime}}\delta_{c^{\prime}f^{\prime}}\,\operatorname{Tr}\big(S^{d}_{ad^{\prime}}(x,0)\gamma^{\rho}\tilde{S}^{u}_{be^{\prime}}(x,0)\gamma^{\mu}\big)\operatorname{Tr}\big(S^{d}_{cc^{\prime}}(x,0)\gamma^{5}\tilde{S}^{d}_{ff^{\prime}}(x,0)\gamma^{5}\big)$
$\qquad\times\operatorname{Tr}\big(S^{d}_{da^{\prime}}(x,0)\gamma^{\sigma}\tilde{S}^{u}_{eb^{\prime}}(x,0)\gamma^{\nu}\big)$
$+$ $\delta_{ae}\delta_{bd}\delta_{cf}\,\delta_{a^{\prime}d^{\prime}}\delta_{b^{\prime}e^{\prime}}\delta_{c^{\prime}f^{\prime}}\,\operatorname{Tr}\big(S^{d}_{ad^{\prime}}(x,0)\gamma^{\rho}\tilde{S}^{u}_{be^{\prime}}(x,0)\gamma^{\mu}\big)\operatorname{Tr}\big(S^{d}_{cc^{\prime}}(x,0)\gamma^{5}\tilde{S}^{d}_{ff^{\prime}}(x,0)\gamma^{5}\big)$
$\qquad\times\operatorname{Tr}\big(S^{d}_{da^{\prime}}(x,0)\gamma^{\sigma}\tilde{S}^{u}_{eb^{\prime}}(x,0)\gamma^{\nu}\big)$
- $\delta_{ad}\delta_{be}\delta_{cf}\,\delta_{a^{\prime}d^{\prime}}\delta_{b^{\prime}e^{\prime}}\delta_{c^{\prime}f^{\prime}}\,\operatorname{Tr}\big(S^{d}_{ad^{\prime}}(x,0)\gamma^{\rho}\tilde{S}^{u}_{be^{\prime}}(x,0)\gamma^{\mu}\big)\operatorname{Tr}\big(S^{d}_{cc^{\prime}}(x,0)\gamma^{5}\tilde{S}^{d}_{ff^{\prime}}(x,0)\gamma^{5}\big)$
$\qquad\times\operatorname{Tr}\big(S^{d}_{da^{\prime}}(x,0)\gamma^{\sigma}\tilde{S}^{u}_{eb^{\prime}}(x,0)\gamma^{\nu}\big)$
- $\delta_{af}\delta_{be}\delta_{cd}\,\delta_{a^{\prime}e^{\prime}}\delta_{b^{\prime}f^{\prime}}\delta_{c^{\prime}d^{\prime}}\,\operatorname{Tr}\big(S^{d}_{ad^{\prime}}(x,0)\gamma^{\rho}\tilde{S}^{u}_{be^{\prime}}(x,0)\gamma^{\mu}\big)\operatorname{Tr}\big(S^{d}_{cc^{\prime}}(x,0)\gamma^{5}\tilde{S}^{d}_{ff^{\prime}}(x,0)\gamma^{5}\big)$
$\qquad\times\operatorname{Tr}\big(S^{d}_{da^{\prime}}(x,0)\gamma^{\sigma}\tilde{S}^{u}_{eb^{\prime}}(x,0)\gamma^{\nu}\big)$
$+$ $\delta_{ae}\delta_{bf}\delta_{cd}\,\delta_{a^{\prime}e^{\prime}}\delta_{b^{\prime}d^{\prime}}\delta_{c^{\prime}f^{\prime}}\,\operatorname{Tr}\big(S^{d}_{ad^{\prime}}(x,0)\gamma^{\rho}\tilde{S}^{u}_{be^{\prime}}(x,0)\gamma^{\mu}\big)\operatorname{Tr}\big(S^{d}_{cc^{\prime}}(x,0)\gamma^{5}\tilde{S}^{d}_{ff^{\prime}}(x,0)\gamma^{5}\big)$
$\qquad\times\operatorname{Tr}\big(S^{d}_{da^{\prime}}(x,0)\gamma^{\sigma}\tilde{S}^{u}_{eb^{\prime}}(x,0)\gamma^{\nu}\big)$
- $\delta_{ad}\delta_{bf}\delta_{ce}\,\delta_{a^{\prime}f^{\prime}}\delta_{b^{\prime}e^{\prime}}\delta_{c^{\prime}d^{\prime}}\,\operatorname{Tr}\big(S^{d}_{ad^{\prime}}(x,0)\gamma^{\rho}\tilde{S}^{u}_{be^{\prime}}(x,0)\gamma^{\mu}\big)\operatorname{Tr}\big(S^{d}_{cc^{\prime}}(x,0)\gamma^{5}\tilde{S}^{d}_{ff^{\prime}}(x,0)\gamma^{5}\big)$
$\qquad\times\operatorname{Tr}\big(S^{d}_{da^{\prime}}(x,0)\gamma^{\sigma}\tilde{S}^{u}_{eb^{\prime}}(x,0)\gamma^{\nu}\big)$
$+$ $\delta_{af}\delta_{bd}\delta_{ce}\,\delta_{a^{\prime}f^{\prime}}\delta_{b^{\prime}d^{\prime}}\delta_{c^{\prime}e^{\prime}}\,\operatorname{Tr}\big(S^{d}_{ad^{\prime}}(x,0)\gamma^{\rho}\tilde{S}^{u}_{be^{\prime}}(x,0)\gamma^{\mu}\big)\operatorname{Tr}\big(S^{d}_{cc^{\prime}}(x,0)\gamma^{5}\tilde{S}^{d}_{ff^{\prime}}(x,0)\gamma^{5}\big)$
$\qquad\times\operatorname{Tr}\big(S^{d}_{da^{\prime}}(x,0)\gamma^{\sigma}\tilde{S}^{u}_{eb^{\prime}}(x,0)\gamma^{\nu}\big)\Big)+\text{27638 similar terms}.$
Here, $\widetilde{S}(x)=C\,S^{T}(x)\,C$ and $S^{u(d)}(x)$ are the $u$ and $d$ quarks’ propagators, with color indices $a$, $b$, $\ldots$. The color deltas ensure the correct color contractions in the traces. The next step in the QCD side involves employing the light-quark propagator in position space:
[…]
$+\frac{m_{q}g_{s}}{32\pi^{2}}G^{\mu\nu}_{ab}\sigma_{\mu\nu}\left[\ln\left(\frac{-x^{2}\Lambda^{2}}{4}\right)+2\gamma_{E}\right]+\cdots.$
In these expressions, $q=u,d$, $\gamma_{E}\approx 0.577$ denotes the Euler–Mascheroni constant, $\Lambda$ is a scale parameter, $G^{2}\equiv G^{A}_{\mu\nu}G^{\mu\nu}_{A}$, $G_{\mu\nu}^{ab}\equiv G^{\mu\nu}_{A}\,t^{A}_{ab}$, $A=1,2,\ldots,8$, and $t^{A}=\lambda^{A}/2$, with $\lambda^{A}$ being the Gell–Mann matrices. The terms in Eq. (10) include the perturbative part, represented by the first term corresponding to the free propagator of a light quark, emission of a single gluon and the non-perturbative contributions, which involve other gluonic effects such as the two-gluon condensate. For the two–gluon condensate, $\langle 0|G^{A}_{\alpha\beta}(x)G^{B}_{\alpha^{\prime}\beta^{\prime}}(0)|0\rangle$, which leads to the four–dimension non–perturbative contribution, we consider the first term of the Taylor expansion for the gluon field at $x=0$. We utilize Barsbay:2022gtu :
$\langle 0|G_{\alpha\beta}^{A}(0)\,G_{\alpha^{\prime}\beta^{\prime}}^{B}(0)|0\rangle=\frac{\langle G^{2}\rangle}{96}\,\delta^{AB}\left(g_{\alpha\alpha^{\prime}}g_{\beta\beta^{\prime}}-g_{\alpha\beta^{\prime}}g_{\alpha^{\prime}\beta}\right),$ (11)
together with the SU(3) color identity:
$t^{a}_{b}t^{a^{\prime}}_{b^{\prime}}=\frac{1}{2}(\delta_{ab^{\prime}}\delta_{a^{\prime}b}-\frac{1}{3}\delta_{ab}\delta_{a^{\prime}b^{\prime}}).$ (12)
In fact, the correlator $\Pi^{\mathrm{Phys}}(p^{2})$ can be rewritten using the dispersion relation:
$\Pi^{\text{Phys}}(p^{2})=\int_{M^{\prime 2}}^{\infty}\frac{\rho^{\text{Phys}}(s)}{s-p^{2}}ds+\cdots,$ (13)
with $M^{\prime}=(4m_{d}+2m_{u})$. The spectral density is
$\rho^{\text{Phys}}(s)=f_{\Delta^{0}\Delta^{0}}^{2}\,\delta(s-m_{\Delta^{0}\Delta^{0}}^{2})+\rho^{h}(s)\,\theta(s-s_{0}),$ (14)
where $\theta(z)$ is the step function and $s_{0}$ is the continuum threshold. Higher resonances and continuum contributions to $\rho^{\mathrm{Phys}}(s)$ are represented through the unknown hadronic spectral density $\rho^{h}(s)$. It is evident that $\rho^{\mathrm{Phys}}(s)$ yields the following expression:
$\Pi^{\text{Phys}}(p^{2})=\frac{f_{\Delta^{0}\Delta^{0}}^{2}}{m_{\Delta^{0}\Delta^{0}}-p^{2}}+\int_{s_{0}}^{\infty}\frac{\rho^{h}(s)}{s-p^{2}}ds.$ (15)
The theoretical calculation of $\Pi^{\mathrm{OPE}}(p^{2})$ is performed in the deep Euclidean region ($p^{2}\ll 0$) via OPE. By analytically continuing $\Pi^{\mathrm{OPE}}(p^{2})$ to the Minkowski domain and taking its imaginary part, the two–point spectral density $\rho^{\mathrm{OPE}}(s)$ is obtained, as presented below. In the deep Euclidean region ($p^{2}\ll 0$), we perform the Borel transformation $\mathcal{B}$ to eliminate subtraction terms in the dispersion relation and to suppress the contributions from higher resonances and the continuum. Applying the Borel transform $\mathcal{B}$ in this domain yields:
$\mathcal{B}\Pi^{\text{Phys}}(p^{2})=f_{\Delta^{0}\Delta^{0}}^{2}\,e^{-m_{\Delta^{0}\Delta^{0}}^{2}/M^{2}}+\int_{s_{0}}^{\infty}\rho^{h}(s)\,e^{-s/M^{2}}\,ds,$ (16)
where $M^{2}$ denotes the Borel parameter and $s_{0}$ is the continuum threshold. To extract physical information from the correlation function, after applying the Borel transformation to both sides of the sum rule, the invariant amplitudes $\Pi^{\mathrm{Phys}}(p^{2})$ and $\Pi^{\mathrm{QCD}}(p^{2})$ are equated. Following this procedure, assuming hadron–parton duality, $\rho^{h}(s)\simeq\rho^{\mathrm{QCD}}(s)$ in the duality region, we subtract the continuum part from the QCD side, yielding:
$f_{\Delta^{0}\Delta^{0}}^{2}e^{-m_{\Delta^{0}\Delta^{0}}^{2}/M^{2}}=\Pi^{\text{QCD}}(s_{0},M^{2}),$ (17)
where
$\Pi^{\text{QCD}}(s_{0},M^{2})=\int_{M^{\prime 2}}^{s_{0}}ds\,\rho(s)\,\ e^{-s/M^{2}}.$ (18)
The spectral density $\rho(s)$ corresponds to the imaginary part of $\Pi^{\mathrm{OPE}}(p^{2})$ and is given by $\rho(s)=\frac{1}{\pi}\mathrm{Im}[\Pi(s)]$. Under the assumptions $m_{u}\to 0$ and $m_{d}\to 0$, the explicit form of $\rho(s)$ reads:
$\rho(s)=$ $-\frac{47\,s^{7}}{6242697216000\,\pi^{10}}+\frac{527\,\langle G^{2}\rangle\,s^{5}\,g_{s}^{2}}{1651507200\,\pi^{10}}+\frac{s^{4}\,\langle\bar{q}q\rangle^{2}\,g_{s}^{2}}{376233984\,\pi^{8}}-\frac{1961\,s^{4}\,\langle\bar{q}q\rangle^{2}}{209018880\,\pi^{6}}-\frac{331\,s^{4}\,\langle\bar{q}q\rangle^{2}}{52254720\,\pi^{6}}$
$-\frac{41\,s^{4}\,\langle\bar{q}q\rangle^{2}\,g_{s}^{2}}{6449725440\,\pi^{8}}-\frac{23\,s^{4}\,\langle\bar{q}q\rangle^{2}}{20901888\,\pi^{6}}+\frac{679\,s^{3}\,\langle\bar{q}q\rangle^{2}\,m_{0}^{2}}{7464960\,\pi^{6}}+\frac{13\,s^{3}\,\langle\bar{q}q\rangle^{2}\,m_{0}^{2}}{207360\,\pi^{6}}+\frac{17\,s^{3}\,m_{0}^{2}\,\langle\bar{q}q\rangle^{2}}{1492992\,\pi^{6}}.$ (19)
In the following, using the shorthand
$\Pi^{\prime}(s_{0},M^{2})=\frac{d}{d(-1/M^{2})}\,\Pi(s_{0},M^{2}),$ (20)
and after straightforward algebra, we obtain the mass and the residue
$m^{2}_{\Delta^{0}\Delta^{0}}=\frac{\Pi^{\prime}(s_{0},M^{2})}{\Pi(s_{0},M^{2})}=\frac{\displaystyle\int_{M^{\prime 2}}^{s_{0}}ds\,s\,\rho(s)\,e^{-s/M^{2}}}{\displaystyle\int_{M^{\prime 2}}^{s_{0}}ds\,\rho(s)\,e^{-s/M^{2}}},$ (21)
and
$f_{\Delta^{0}\Delta^{0}}^{2}=e^{m_{\Delta^{0}\Delta^{0}}^{2}/M^{2}}\,\Pi(s_{0},M^{2}).$ (22)
III Numerical analysis
For the numerical evaluation of the obtained expressions for the mass and residue of the $\Delta^{0}\Delta^{0}$ dibaryon states with $J^{P}=2^{+}$, a set of input parameters, including the gluon and quark condensates, is required. These parameters are listed in Table I.
$\langle\bar{q}q\rangle$   $(0.24\pm 0.01)^{3}~\text{GeV}^{3}$ Belyaev:1982sa   $m_{0}^{2}$   $(0.8\pm 0.1)~\text{GeV}^{2}$Belyaev:1982sa   $\langle\alpha_{s}G^{2}\rangle$   $(0.012\pm 0.004)~\text{GeV}^{4}$ Belyaev:1982cd
$\langle\alpha_{s}\rangle$   $(0.118\pm 0.005)$ Furstenau:1994ra   $\langle g_{s}^{2}\rangle$   $4\pi\alpha_{s}$
Table 1: Key input parameters and their values used in the analysis.
At this point in the analysis, it is crucial to determine the appropriate working windows for the two auxiliary parameters involved in QCD sum rule calculations of physical observables, namely the Borel parameter $M^{2}$ and the continuum threshold $s_{0}$. While these parameters are not physical in themselves, they are subject to specific constraints within the QCD sum rule framework and cannot be selected arbitrarily.
A key requirement is that physical quantities such as masses and residue should remain largely stable within certain ranges of $M^{2}$ and $s_{0}$. Identifying these stability regions is essential for ensuring the reliability of the results, as it helps minimize theoretical uncertainties and validates the consistency of the QCD sum rule approach.
The parameters must be selected to ensure that the pole contribution (PC) dominates in the physical observables extracted via the sum rule method. Additionally, the stability of these results with respect to $M^{2}$ and the convergence of the OPE are critical constraints in the numerical analysis. These conditions can be quantitatively assessed by introducing the following expressions:
$\text{PC}=\frac{\Pi(M^{2},s_{0})}{\Pi(M^{2},\infty)},$ (23)
and
$R(M^{2})=\frac{\Pi^{\text{Dim}N}(M^{2},s_{0})}{\Pi(M^{2},s_{0})}.$ (24)
Here, we impose the following criteria: The upper bound of the Borel parameter $M^{2}$ is determined by ensuring that the pole contribution is considerably large compared to each of the higher excited states and continuum. In technical language, we require that the pole contribution be as maximum as possible in the exotic dibaryon channel: We set $\mathrm{PC}\geq 0.2$. Moreover, the lower bound of $M^{2}$ is determined by the requirement of convergence of the OPE. This criterion implies that the perturbative contribution must dominate over the non-perturbative terms, and the effects of higher-dimensional operators should remain suppressed. To ensure this, we require that the combined contribution of the last three non-perturbative terms does not exceed $5\%$ of the total (perturbative plus non-perturbative) contributions, i.e., $R(M^{2})\leq 0.05$. Moreover, the chosen range for $s_{0}$ should lie significantly above the production threshold of two $\Delta^{0}$ baryons, which is approximately $4m_{\Delta^{0}}^{2}\simeq 6.07~\mathrm{GeV}^{2}$. The range of $s_{0}$ may be selected as the minimal value that ensures a stable and reliable Borel window. Alternatively, $s_{0}$ can be fixed by fulfilling additional constraints so as to maximize the PC. Note that for multiquark hadrons, or more precisely exotic states, the available information is often scarce. In such situations, $s_{0}$ can be estimated based on the aforementioned constraints.
We then choose the parameters $(2m_{\Delta^{0}\Delta^{0}}+0.2)^{2}~\mathrm{GeV}^{2}\leq s_{0}\leq(2m_{\Delta^{0}\Delta^{0}}+0.4)^{2}~\mathrm{GeV}^{2}$, which define an energetically feasible window for exciting the $\Delta^{0}\Delta^{0}$ system to its first excited state. In addition, the Borel parameter is chosen as $8.0~\mathrm{GeV}^{2}\leq M_{B}^{2}\leq 10.0~\mathrm{GeV}^{2}$, which ensures the necessary conditions for a reliable QCD sum rule analysis.
Figures 1 and 2 present the mass and decay constant of the $\Delta^{0}\Delta^{0}$ dibaryon as functions of $M^{2}$ and $s_{0}$. Figure 1 (left panel) demonstrates that the mass remains relatively stable across the chosen range of the Borel parameter $M^{2}$ and shows very good stability. In the right panel of Figure 1, a mild dependence of the mass on the continuum threshold $s_{0}$ is observed. As shown in Figure 2, though the residue shows a good stability with respect to the variations of $M^{2}$ in its working region, it indicates a residual sensitivity to variations in $s_{0}$. As a result, the requirements of the method are satisfied.
Figure 1: The left panel shows the variation of the $\Delta^{0}\Delta^{0}$ dibaryon mass with respect to $M^{2}$ at fixed $s_{0}$, while the right panel illustrates its dependence on $s_{0}$ at fixed $M^{2}$.
Figure 2: The left panel shows the variation of the $\Delta^{0}\Delta^{0}$ dibaryon mass with respect to $M^{2}$ at fixed $s_{0}$, while the right panel illustrates its dependence on $s_{0}$ at fixed $M^{2}$.
To extract the mass $m_{\Delta^{0}\Delta^{0}}$ and residue $f_{\Delta^{0}\Delta^{0}}$ of the $\Delta^{0}\Delta^{0}$ dibaryon with $J^{P}=2^{+}$, we perform calculations using the range of values for the Borel parameter $M^{2}$ and the continuum threshold $s_{0}$. The final values are obtained by averaging over the working intervals defined above. The results are as follows:
$m_{\Delta^{0}\Delta^{0}}=2426^{+101}_{-108}~\text{MeV},\quad f_{\Delta^{0}\Delta^{0}}=\left(2.30^{+0.49}_{-0.45}\right)\times 10^{-4}~\text{GeV}^{8}$ (25)
The uncertainties in the mass and decay residue arise from the variation between the extreme values of $m_{\Delta^{0}\Delta^{0}}$ and $f_{\Delta^{0}\Delta^{0}}$. The central mass value lies approximately $38~\mathrm{MeV}$ below the $2m_{\Delta^{0}}$ threshold, indicating a possible bound molecular state of the scalar $\Delta^{0}\Delta^{0}$ dibaryon. Compared to the calculated $\Delta^{0}\Delta^{0}$ dibaryon in the $J^{P}=0^{+}$ channel, where a mass of $m_{\Delta^{0}\Delta^{0}}=2326^{+114}_{-126}~\mathrm{MeV}$ has been reported Mutuk:2022zgn , the mass obtained in the $J^{P}=2^{+}$ channel is higher by about $100~\mathrm{MeV}$, leading to a weaker binding. It is also worth noting that the $\Omega\Omega$ dibaryon, an exotic particle composed entirely of strange quarks, has been investigated in both the $J^{P}=0^{+}$ and $J^{P}=2^{+}$ channels Chen:2019vdh . In contrast to the $\Delta^{0}\Delta^{0}$ dibaryon, the predicted mass of the $\Omega\Omega$ state in the tensor channel is lower than that in the scalar channel, indicating a stronger binding for the tensor state.
It is important to emphasize that the present analysis excludes the influence of radiative corrections. Although the mass, extracted from the ratio of two sum rules, is relatively insensitive to such corrections, the residue calculated from an individual sum rule can exhibit considerable fluctuations. Furthermore, next-to-leading order (NLO) perturbative corrections to baryon sum rules have been reported as significant Krasnikov:1982ea ; Groote:1999zp . Similarly, while NLO $\alpha_{s}$ corrections to the perturbative component in the OPE for light tetraquark currents are sizable, their overall impact on the sum rules remains limited owing to the predominance of nonperturbative condensate contributions Groote:2014pva . Furthermore, given that the analysis of the $\Delta^{0}\Delta^{0}$ dibaryon with $J^{P}=2^{+}$ suggests the existence of a bound state, the estimation of its size can be interesting. Based on the definition of binding energy, the size of the $\Delta^{0}\Delta^{0}$ dibaryon can be determined using the following expression:
$r\sim\frac{1}{\sqrt{2\mu\,B_{E}}}$ (26)
where the binding energy is $B_{E}=2m_{\Delta^{0}}-m_{\Delta^{0}\Delta^{0}}$. In this expression, $r$ represents the distance between the components, while $\mu=\frac{m_{1}m_{2}}{m_{1}+m_{2}}$ corresponds to the reduced mass of the two-hadron system. The size of the tensor state is found to be approximately $r\simeq 0.90~\mathrm{fm}$, nearly twice that of the $\Delta^{0}\Delta^{0}$ dibaryon in the $J^{P}=0^{+}$ state. As is clear, both the tensor and scalar states remain below the confinement radius of 1 fm  Guo:2017jvc . In the case of the tensor state, while it may suggest a molecular configuration, the radius indicates an overlap between two $\Delta^{0}$ baryons, making the molecular description questionable. The obtained size suggests a compact hexaquark structure for this particle. However, calculating other properties of this state may help in achieving a better understanding of its internal configuration. In contrast to the $\Delta^{0}\Delta^{0}$ dibaryon, which represents a light-quark exotic dibaryon, it is of particular interest to estimate the radius of the $\Omega\Omega$ dibaryon mentioned above. Based on the binding energy calculated in Ref. Chen:2019vdh , and assuming the same quantum state as considered in our study, the resulting radii are found to be approximately $r\simeq 1.24~\mathrm{fm}$ for the $J^{P}=0^{+}$ state and $r\simeq 0.97~\mathrm{fm}$ for the $J^{P}=2^{+}$ state. These results indicate that the tensor state remains within the typical confinement radius.
IV CONCLUDING NOTES
The investigation of newly observed exotic hadronic states constitutes a primary focus within contemporary hadron physics research. In this study, we explore tensor $\Delta^{0}\Delta^{0}$ dibaryon states with spin-parity $J^{P}=2^{+}$ using the framework of QCD sum rules. We construct interpolating currents for the tensor $\Delta^{0}\Delta^{0}$ dibaryon, and calculate the corresponding spectral densities and two-point correlation functions, including contributions from condensates up to dimension nine. The numerical analysis indicates that the tensor current leads to stable mass sum rules. Based on this, we provide a reliable mass prediction for the tensor $\Delta^{0}\Delta^{0}$ dibaryon with $J^{P}=2^{+}$, obtaining $m_{\Delta^{0}\Delta^{0},\,2^{+}}=2426^{+101}_{-108}~\mathrm{MeV}.$ This result points to the possible existence of a bound state, with an estimated binding energy of approximately $38~\mathrm{MeV}$. The residue was determined to be $f_{\Delta^{0}\Delta^{0}}=\left(2.30^{+0.49}_{-0.45}\right)\times 10^{-4}~\mathrm{GeV}^{8}.$ Additionally, we estimated the spatial size of the tensor $\Delta^{0}\Delta^{0}$ dibaryon to be approximately $r\simeq 0.90~\mathrm{fm}$. The calculated binding energy suggests that the tensor $\Delta^{0}\Delta^{0}$ dibaryon is a potentially stable system. We hope that these findings will contribute to future theoretical investigations and experimental searches for dibaryon states.